if A(3,4) and B is a variable point on the lines modx=6.If AB<=4 then what is the no. of position of B with integral co-ordinates?

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Sneha,

Ans:- Let the point b is (6,k)  { B can not be (-6,k) then the inegulity (1) does not hold}

Then √(3²+(k-4)²<=4..................(1)

or 9+(k-4)²<=16

or  4- √7<=k <=4+ √7

Hence the integral soln are,

k= 2,3,4,5,6

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Sneha!!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR